# Single Image Test-Time Adaptation for Segmentation

### v1

@@ 4.2 COCO $\rightarrow$ VOC
In Appendix E, the results with optimal hyper-parameters for each method, corruption kind, and level are shown. This time, the results between different methods are much smaller. The PL consistently outperforms all other methods at all the corruption levels. Interestingly, the dIoU methods shows much stronger performance than in the GTA5-C experiments.
@@
Again, only Ref, PL, AugCo and Ent are used for further analysis. The relationship between the non-adapted (NA) performance and the performance improvement on individual images for different methods is visualized in Figure 7. The distribution of initial non-adapted $\text{m}\overline{\text{IoU}}_{i}$ is different. The initial model is stronger than the GTA5 model. All methods show similar behavior - more improvement is achieved on images with a lower initial score, Ref and PL significantly outperform the other methods and again, Ref show better performance on images with low initial scores compared to other methods.
@@
The relationship of segmentation prediction entropy and $\text{m}\overline{\text{IoU}}_{i}$ improvement by adaptation is shown in Figure 8, supporting the notion that the entropy of prediction before adaptation is a good proxy for $\text{m}\overline{\text{IoU}}_{i}$.
Figure 9: TTA $\text{m}\overline{\text{IoU}}_{i}$ error reduction (%) of COCO model on the VOC test dataset with hyper-parameters selected on COCO-C.
@@
The results on the VOC test set are shown in Figure 9. PL slightly outperforms the other methods, but all the methods improved over the non-adapted baseline.
@@
5 Conclusions and Discussion
The first finding is that all the methods minimizing the CE loss can be improved by optimizing the IoU loss. Optimizing all the network parameters, compared to optimizing the normalization parameters only, is shown to bring large performance gains for the Ref method, making it another important hyper-parameter to consider. With these improvements, we identify PL and Ref as the overall best-performing methods. The performance is improved by 3.51 % and 3.28 % on GTA5-C and COCO-C validation sets, while without IoU loss, the improvements would have been by 1.7 % 2.16 % only. This could be explained by the fact that IoU loss is better suited for highly-imbalanced data. The imbalance is even stronger for the single-image setup.
The experiments further reveal that even though the methods achieve impressive performance gains when optimizing hyper-parameters for each corruption and corruption level separately, there is a huge performance drop when evaluating with a single set of hyper-parameters. This results in negative performance on many corruptions. It is also shown that Ref performs better compared to PL when the segmentation quality before adaptation is low, while it may deteriorate performance on already high-quality masks. Entropy is shown to be a good estimate of the quality. This reveals the great potential of learnt refinement modules for severe domain shifts but also highlights the need for improvements in the training of the module to ensure performance is not degraded for high-quality masks.
The Ent baseline is often reported as too unstable for single-image adaptation, producing negative results. In contrast, [41], and our experiments show positive results. The difference is that our work and [41] do not update the batch normalization statistics like the original work proposes, only the learnable parameters.
Finally, the test results further highlight that it is not possible to find a single set of hyper-parameters for the tested TTA, since the results do not always reflect those on the validation set. If the domain shift kind is known, the validation set can be tuned to these conditions, but that undermines the idea of adapting to an arbitrary unknown domain shift - if the conditions are known in advance, there are many techniques to exploit that information.
Limitations First, a single isolated image adaptation by optimizing model parameters has a high computational cost, as a backward operation is required for each iteration. If such methods are to be broadly adopted, this issue should be addressed. Further, many of the methods evaluated in this work were not designed with single-image adaptation in mind, and modifications were necessary. Poor performance in our benchmarks does not invalidate the performance of the methods in the setup these were proposed for. Finally, some of the methods have not made the code public. The code with the methods reimplementation will be released.
Appendix A Related setups
Domain adaptation methods share a lot in common with semi-supervised learning [31, 33] and may exploit strategies such as generative models [39, 47]. In practice, this is often infeasible since source data may not be available for example for privacy reasons, or we may only have a small number of target domain images available. In continually evolving environments, the distribution may have already changed by the time adaptation on a whole target dataset is completed. Various modifications of the traditional domain adaptation scenario tackling the aforementioned limitations have recently emerged, for example by considering no access to source data or a continual domain shift [22, 41, 44, 3]. In particular, test-time adaptation methods assume no source data is available and aim to exploit the information from as little as a single target domain image.
TTA can also be described as source-free unsupervised domain adaptation. A model trained on a source domain has to adapt to a target domain without access to the source data and without any labels. Let us first describe related terms and also the difference between similar but different setups that may be easily confused:
@@ Appendix A Related setups
Domain generalization aims for a strong model that would generalize to unseen domains without any adaptation. Common approaches are domain-invariant representation learning, data augmentation or data generation. In contrast, this work focuses on adapting a pre-trained model to become a specialist in the current domain. It was shown in [41] that a stronger, more general model can lead to better adaptation results, making these directions complementary. However, without adaptation, training a very general model on a large set of distributions may harm the model’s performance on the individual distributions when compared to specialized models with carefully optimized data augmentation [1, 36].
@@ B.1 Entropy Minimization (Ent)
The Ent method minimizes the entropy of the segmentation predictions. In the context of learning with limited supervision, it was proposed in [11] for semi-supervised learning. In TTA, there is no labelled set that could be leveraged as regularization like in semi-supervised learning but the methods were shown to work for TTA as well [43]. It was also shown that larger batch size and updating the parameters of the normalization-layers only improve stability of the method. But on segmentation, a dense-prediction task, adapting to a single image can lead to positive results, especially when not updating the batch normalization statistics but only the learnable parameters like in [41].
@@
In this work, the batch normalization [14] statistics are not updated since it relies on the presence of batch normalization layers while many recent architectures use other normalization layers such as layer normalization [2].
@@ B.2 Pseudolabelling (PL)
Pseudolabelling also comes from semi-supervised learning [18] and is based on the idea of using the segmentation prediction of an image (prediction for each class binarized through argmax) as ground truth to optimize the model. In effect, it is the same as pseudo-labelling since both methods reduce class overlap. However, pseudolabelling has the advantage of allowing for different loss function. While the CE loss is typically optimized, our experiments show that in the single image setup, IoU leads to superior results.
@@ B.3 Adversarial Transformation Invariance (Adv)
This method is an extension of TIPI (Test-Time Adaptation with Transformation Invariance) by [27] to image segmentation. The main idea is to make the network invariant to adversarial transformation of the input image as a representative of domain shifts.
@@
In forward KL, p corresponds to the model prediction while q to the ground truth. Please note that, as suggested in [27], the reverse KL is used in the proposed method where the input arguments to the function are switched, compared to forward KL. Another important implementation detail is that the gradients should not flow through $\text{s}^{\prime}_{i}$ the tensor needs to be detached before the loss computation.
The same adversarial attacks in terms of the ground truth as for the IoU estimation and mask refinement methods are used to generate $x^{\prime}$ but the computational complexity is reduced by using the Fast Gradient Sign Attack (FGSM) proposed in [10].
Importantly, two sets of batch normalization statistics are kept in [27], which is not done in our work due to the aim for general methods that do nto assuem the presence of specific network layers. This may be the reason why the methods perform poorly in our experiments. Another thing we ntoed is the high variance of the KL loss.
@@ B.4 Augmentation Consistency (AugCo)
In [29], an auxiliary information entropy loss preventing trivial solutions is also optimized. This loss requires running class-frequency statistics and is not applicable to the single-image setup. Further, an adaptive threshold based on per-class confidence distirbution in a batch of images is computed iinstead of a fixed threshold, which is also not applicable in our setup.
@@ Appendix D Synthetic corruptions
The corruptions used in our experiments are a subset of the corruptions from [12]. An overview of the corruptions, as well as implementation details, can be found in Table 4
corruption
description
brightness
is an additive intensity transformation, $x_{c}=\text{clip}(x+b)$ where $b$ controls the level.
contrast
is a multiplicative intensity transformation, $x_{c}=b(x-\overline{x})+x$ where $b$ controls the level and $\overline{x}$ is a per-channel mean of the image intensities.
frost
first crops a portion of one of the frost image templates at a random location of the same size as the input image, $x_{\text{f}}$. Then we compute $x_{c}=b_{1}x+b_{2}x_{\text{f}}$ where the weights $b_{1},b_{2}$ control the level.
fog
first generates a heightmap $x_{\text{h}}$ using the diamond-square algorithm [7], where the wibble is controlled by a parameter $b_{1}$. It is then combined with the input image as $x_{c}=\frac{(x+b_{1}\cdot x_{\text{h}})\cdot\max(x)}{\max(x)+b_{1}}$.
gaussian noise
is generated as $x_{c}=x+n$ where $n\sim\mathcal{N}(0,b)$ and $b$ controls the level.
shot noise
is generated as $x_{c}\sim\frac{\text{Pois}(x\cdot b,\lambda=1)}{b}$ where Pois denotes the Poisson distribution and $b$ controls the level.
spatter
simulates mud or water spoiling. The main idea of the algorithm is a combination of thresholding and blurring random noise.
defocus blur
first generates a disk kernel K with radius $b_{1}$ and alias blur $b_{2}$. The kernel is then used to filter each of the channels $x_{c}=\text{K}(x)$.
gaussian blur
corrupts the image by gaussian blurring $x_{c}=\mathcal{N}(x,b)$ where $b$ controls the level.
jpeg
is computed as $x_{c}=\text{jpeg}(x,b)$ where jpeg performs the JPEG compression with quality $b$.
Table 4: Corruptions and their implementation details, a subset from [12]. The input of the transformation is an image $x$ normalized to the $(0,1)$ range, the output is a corrupted image $x_{c}$. The clip function limits the values to the $[0,1]$ range. This function is always applied to the output image after the transformation to obtain the final output $x_{f}=\text{clip}(x_{c})$. For more details on the transformations and the values defining the level, please refer to the codebase.
@@ Appendix E Additional experimental results
method
dataset metric NA Ref PL Ent AugCo
Cityscapes $\text{m}\overline{\text{IoU}}_{i}$ 34.40 34.71 37.14 35.11 35.45
$\text{m}\overline{\text{IoU}}_{c}$ 28.71 28.64 30.70 29.09 29.48
ACDC-fog $\text{m}\overline{\text{IoU}}_{i}$ 32.03 35.98 35.67 33.93 31.82
$\text{m}\overline{\text{IoU}}_{c}$ 24.87 27.29 27.52 26.00 24.69
ACDC-night $\text{m}\overline{\text{IoU}}_{i}$ 13.60 14.12 15.09 14.13 14.15
$\text{m}\overline{\text{IoU}}_{c}$ 10.77 10.96 11.53 10.68 11.01
ACDC-rain $\text{m}\overline{\text{IoU}}_{i}$ 33.52 35.61 37.17 35.05 34.36
$\text{m}\overline{\text{IoU}}_{c}$ 26.15 27.40 28.47 26.89 26.66
ACDC-snow $\text{m}\overline{\text{IoU}}_{i}$ 31.54 35.60 34.15 31.89 31.81
$\text{m}\overline{\text{IoU}}_{c}$ 25.28 28.09 27.17 25.39 25.45
method
metric NA Ref PL Ent AugCo
$\text{m}\overline{\text{IoU}}_{i}$ 77.03 78.21 79.83 78.66 77.75
$\text{m}\overline{\text{IoU}}_{c}$ 63.31 64.54 67.13 65.02 63.27
Table 6: VOC test dataset results. Hyper-parameters were selected for overall best performance on COCO-C. Best results for each dataset and metric are highlighted
@@
Figures 16 (GTA5-C) and 15 (COCO-C) show the results when optimal hyper-parameters are determined for each corruption and level separately. These results show that the performance improves substantially compared to the scenario with overall optimal hyper-parameters for all corruptions and levels. Further, while on the GTA5-C dataset, Ref sinigificantly outperforms the other methods, PL is better in the VOC-C benchmark.
Figure 15: Comparison of total error reduction across methods per corruption levels on COCO-C dataset
Figure 16: Comparison of total error reduction across methods per corruption levels on GTA5-C dataset
@@
The numerical results on the Cityscapes and ACDC test dataset can be found in Table 5, showing different methods may perform better depending on the evaluation metric. The same results for the VOC test dataset are shown in Table 6.
Additional experiments The refinement module is trained to predict a clean mask based on a corrupted mask simulating masks processed by the model under domain shift. The clean mask can be the segmentation prediction on clean, non-corrupted images or, when available, ground truth masks can be used instead. Comparison of these two choices is shown in Table 7. The results are somewhat inconclusive - for the COCO model it can be seen that the model trained on predictions is substantially better than the one trained on ground truth. The GTA5 model performs similarly in both cases. One could argue that learning with GT can compensate for some of the mistakes even source distribution images, since prediction form output space back to output space is different then prediction from image to output space. On the other hand, when predictions differ significantly from ground truth even on source distirbution images, it can result in noiser data and more difficult training. The choice should be validated experimentally for each model and dataset.

### v2

@@ 5.3 Experiment results
In Figure 9, the results with optimal hyper-parameters for each method, corruption kind and level are shown. The results again improve substantially but this time, the differences between methods are smaller. Interestingly, the dIoU method performs much stronger than in the GTA5-C experiments.
Figure 10(b) compares the overall method performance on the SITTA training set. An oracle option is introduced where the method with the best results is picked for each image. There is a significant gap between the oracle and other methods, which further highlights that different methods are good in various cases and understanding the strengths of each methods can lead to significantly improved performance.
Figure 9: COCO-C error reduction difference (%) between overall optimal hyperparameters and hyper-parameters selected for each corruption kind separately. The hyper-parameters were selected on the COCO-C.
(a) Comparison of the overall performance of different methods on the GTA5-C validation set.
(b) Comparison of the overall performance of different methods on the COCO-C validation set.
Figure 10: The difference between non-adapted (NA) $\text{m}\overline{\text{IoU}}_{i}$
@@
Again, only Ref, PL, AugCo, and Ent are used for further analysis. The relationship between the non-adapted (NA) performance and the performance improvement on individual images for different methods is visualized in Subfigure 8(a). The distribution of initial non-adapted $\text{m}\overline{\text{IoU}}_{i}$ is different. The initial model is stronger than the GTA5 model. All methods show similar behavior - more improvement is achieved on images with a lower initial score.
@@
The relationship of segmentation prediction entropy and $\text{m}\overline{\text{IoU}}_{i}$ improvement by adaptation is shown in Subfigure 8(b), supporting the notion that the entropy of prediction before adaptation is a good proxy for $\text{m}\overline{\text{IoU}}_{i}$.
method
metric NA Ref PL Ent AugCo
$m\overline{\mathrm{IoU}}_{i}$ 77.00 77.62 79.67 78.85 77.75
$m\overline{\mathrm{IoU}}_{c}$ 63.38 63.38 66.90 65.38 63.10
mIoU 77.20 77.78 77.55 77.01 76.01
mDice 86.16 86.66 86.47 86.05 85.54
Accuracy 94.69 95.18 94.68 94.69 94.76
Table 4: VOC test dataset results. Hyper-parameters were selected for overall best performance on COCO-C. Best results for each dataset and metric are highlighted.
@@
The results on the VOC test set are shown in Figure 6(b) and Table 4. PL outperforms other methods in the proposed metrics, all the methods improved over the non-adapted baseline. When considering standard metrics which do not consider per-image difficulty, Ref is best.
@@
6 Conclusions and limitations
This work investigated the performance of six Single Image Test-Time Adaptation (SITTA) methods on semantic segmentation problems. Two established baselines, two adapted state-of-the-art methods from image classification and continual TTA, and two novel methods were considered. We designed a framework that allows for hyper-parameter tuning and analysis of performance under diverse conditions on synthetic data inspired by Hendrycks & Dietterich (2019) that can be derived from an arbitrary training dataset. We evaluated the methods on real-world datasets.
Experiments on driving-scene datasets dominate the segmentation TTA literature. We followed this practice and experimented with models pretrained on the synthetic GTA5 dataset, evaluating on real-world driving scenes, including datasets with adverse weather conditions such as rain and fog. The common practice is to tune TTA hyperparameters on the Synthia Ros et al. (2016) datasets, however, this dataset contains the same weather and day-night conditions as the test datasets. Also, it is limited to driving datasets with the same set of classes. For this reason, we tuned hyper-parameters on a synthetic dataset derived from the training dataset by applying different corruptions, which gives us control over the domain shift conditions, facilitates analysis and is applicable to any existing dataset. We also added a novel benchmark where we evaluated a COCO-pretrained model on the VOC dataset, focusing on common objects.
We proposed two new methods inspired by ideas from unsupervised domain adaptation and medical imaging TTA. The stronger of the methods based on a learnt refinement module performed best on multiple of the test datasets and we showed that it is powerful on the images with the worst segmentation performance, as measured by IoU. We also showed that if we replace the IoU performance with the entropy of the predicted segmentation mask, which does not require the ground truth to be known, the same behaviour can be observed. This could be used to choose an appropriate method/hyper-parameters in future work.
We explored the effect of previously neglected design choices. Training with a loss function that accounts for class imbalance, a known issue of image segmentation datasets, such as the IoU or dice loss, is standard when training image segmentation models. When considering small batch sizes or even a single image, the class imbalance further increases, but SITTA methods implement baselines such as pseudo-labelling with the CE loss or entropy minimization with equal weight for all pixels. Our results revealed that while SITTA in the standard setting with CE loss did not improve performance much, substituting the CE with IoU improves performance substantially. The performance of pseudo-labelling was improved by 3.51 % and 3.28 % on GTA5-C and COCO-C validation sets while with CE loss, the improvements were by 1.7 % 2.16 % only, respectively. The experiments on whether to update all or normalization parameters only show this design choice significantly impacts results but the right option depends on the settings (method, dataset) - it is an important hyper-parameter to consider. Further, we find that entropy minimization, often reported as unstable for small batch sizes, performs well when the batch-normalization mean and variance are not updated at test time and only the affine parameters of the normalization layers are optimized.
In the GTA5-C synthetic datasets experiments, the refinement SITTA dominates, followed by the pseudo-labelling baseline. While the refinement is significantly better on some of the real-world test datasets, on other ones, pseudo-labelling performs best. In the COCO-C experiments, the top performers swap places: Pseudo-labelling is followed by refinement. On the test dataset, pseudo-labelling remains the best. While the other methods do not perform very well when TTA hyperparameters are optimized across many different corruption kinds, their performance improves when tuning them to specific kinds of domain shift. While there is not a single method performing best over all the test datasets, our results highlight the potential of SITTA for semantic segmentation.
Limitations. To limit the scope of the study, we only focused on adaptation with self-supervised loss functions and no reliance on batch normalization layers. While these methods tend to perform the best, their iterative optimization comes at an increased computational time. Methods alleviating this burden should be explored, such as only adapting to informative samples or methods inspired by efficient model finetuning.
The synthetic validation set created by applying artificial corruptions to the training set may not cover the complexity of real-world domain shifts. Label shift, common in real-world datasets, is not considered - the solutions are typically complementary.
Finally, only two models were considered and the effect of different model architectures on the individual methods is not known. While our work improves the understanding of TTA for semantic segmentation methods, a benchmark for fair and thorough evaluation of the methods is still missing.
Acknowledgements
We greatly appreciate the help of Tomas Jenicek and Ondrej Tybl with proofreading of the manuscript.
Klara Janouskova was supported by the SGS23/173/OHK3/3T/13 research grant, and Jiri Matas was supported by the Technology Agency of the Czech Republic, project No. SS73020004 and project FACIS No. VJ02010041. Tamir Shor and Chaim Baskin were supported by NOFAR MAGNET number 78732 by the Israel Innovation Authority.
Appendix A Related Domain Adaptation Scenarios
@@
Domain generalization aims for a strong model that would generalize to unseen domains without any adaptation. Common approaches are domain-invariant representation learning, data augmentation or data generation. In contrast, this work focuses on adapting a pre-trained model to become a specialist in the current domain. It was shown in Volpi et al. (2022) that a stronger, more general model can lead to better adaptation results, making these directions complementary. However, without adaptation, training a very general model on a large set of distributions may harm the model’s performance on the individual distributions when compared to specialized models with carefully optimized data augmentation Aquino et al. (2017); Steiner et al. (2021).
@@ B.1 Entropy Minimization (Ent)
The Ent method minimizes the entropy of the segmentation predictions. In the context of learning with limited supervision, it was proposed in Grandvalet & Bengio (2004) for semi-supervised learning. In TTA, there is no labelled set that could be leveraged as regularization like in semi-supervised learning but the methods were shown to work for TTA as well Wang et al. (2020b). It was also shown that larger batch size and updating the parameters of the normalization-layers only improve stability of the method. But on segmentation, a dense-prediction task, adapting to a single image can lead to positive results, especially when not updating the batch normalization statistics but only the learnable parameters like in Volpi et al. (2022).
@@
In this work, the batch normalization Ioffe & Szegedy (2015b) statistics are not updated since it relies on the presence of batch normalization layers while many recent architectures use other normalization layers such as layer normalization Ba et al. (2016).
@@ B.2 Pseudolabelling (PL)
Pseudolabelling also comes from semi-supervised learning Lee et al. (2013) and is based on the idea of using the segmentation prediction of an image (prediction for each class binarized through argmax) as ground truth to optimize the model. In effect, it is the same as pseudo-labelling since both methods reduce class overlap. However, pseudolabelling has the advantage of allowing for different loss function. While the CE loss is typically optimized, our experiments show that in the single image setup, IoU leads to superior results.
@@ B.3 Adversarial Transformation Invariance (Adv)
This method is an extension of TIPI (Test-Time Adaptation with Transformation Invariance) by Nguyen et al. (b) to image segmentation. The main idea is to make the network invariant to adversarial transformation of the input image as a representative of domain shifts.
@@
In forward KL, p corresponds to the model prediction while q to the ground truth. Please note that, as suggested in Nguyen et al. (b), the reverse KL is used in the proposed method where the input arguments to the function are switched, compared to forward KL. Another important implementation detail is that the gradients should not flow through $\text{s}^{\prime}_{i}$ the tensor needs to be detached before the loss computation.
The same adversarial attacks in terms of the ground truth as for the IoU estimation and mask refinement methods are used to generate $x^{\prime}$ but the computational complexity is reduced by using the Fast Gradient Sign Attack (FGSM) proposed in Goodfellow et al. (2014).
Importantly, two sets of batch normalization statistics are kept in Nguyen et al. (b), which is not done in our work due to the aim for general methods that do nto assuem the presence of specific network layers. This may be the reason why the methods perform poorly in our experiments. Another thing we ntoed is the high variance of the KL loss.
@@ B.4 Augmentation Consistency (AugCo)
In Prabhu et al. (2021), an auxiliary information entropy loss preventing trivial solutions is also optimized. This loss requires running class-frequency statistics and is not applicable to the single-image setup. Further, an adaptive threshold based on per-class confidence distirbution in a batch of images is computed iinstead of a fixed threshold, which is also not applicable in our setup.
@@ Appendix D Synthetic corruptions
The corruptions used in our experiments are a subset of the corruptions from Hendrycks & Dietterich (2019). An overview of the corruptions, as well as implementation details, can be found in Table 5
corruption description
brightness is an additive intensity transformation, $x_{c}=\text{clip}(x+b)$ where $b$ controls the level.
contrast is a multiplicative intensity transformation, $x_{c}=b(x-\overline{x})+x$ where $b$ controls the level and $\overline{x}$ is a per-channel mean of the image intensities.
frost first crops a portion of one of the frost image templates at a random location of the same size as the input image, $x_{\text{f}}$. Then we compute $x_{c}=b_{1}x+b_{2}x_{\text{f}}$ where the weights $b_{1},b_{2}$ control the level.
fog first generates a heightmap $x_{\text{h}}$ using the diamond-square algorithm Fournier et al. (1982), where the wibble is controlled by a parameter $b_{1}$. It is then combined with the input image as $x_{c}=\frac{(x+b_{1}\cdot x_{\text{h}})\cdot\max(x)}{\max(x)+b_{1}}$.
gaussian noise is generated as $x_{c}=x+n$ where $n\sim\mathcal{N}(0,b)$ and $b$ controls the level.
shot noise is generated as $x_{c}\sim\frac{\text{Pois}(x\cdot b,\lambda=1)}{b}$ where Pois denotes the Poisson distribution and $b$ controls the level.
spatter simulates mud or water spoiling. The main idea of the algorithm is a combination of thresholding and blurring random noise.
defocus blur first generates a disk kernel K with radius $b_{1}$ and alias blur $b_{2}$. The kernel is then used to filter each of the channels $x_{c}=\text{K}(x)$.
gaussian blur corrupts the image by gaussian blurring $x_{c}=\mathcal{N}(x,b)$ where $b$ controls the level.
jpeg is computed as $x_{c}=\text{jpeg}(x,b)$ where jpeg performs the JPEG compression with quality $b$.
Table 5: Corruptions and their implementation details, a subset from Hendrycks & Dietterich (2019). The input of the transformation is an image $x$ normalized to the $(0,1)$ range, the output is a corrupted image $x_{c}$. The clip function limits the values to the $[0,1]$ range. This function is always applied to the output image after the transformation to obtain the final output $x_{f}=\text{clip}(x_{c})$. For more details on the transformations and the values defining the level, please refer to the codebase.
@@ Appendix E Additional experimental results
Additional experiments The refinement module is trained to predict a clean mask based on a corrupted mask simulating masks processed by the model under domain shift. The clean mask can be the segmentation prediction on clean, non-corrupted images or, when available, ground truth masks can be used instead. Comparison of these two choices is shown in Table 6. The results are somewhat inconclusive - for the COCO model it can be seen that the model trained on predictions is substantially better than the one trained on ground truth. The GTA5 model performs similarly in both cases. One could argue that learning with GT can compensate for some of the mistakes even source distribution images, since prediction from output space back to output space is different then prediction from image to output space. On the other hand, when predictions differ significantly from ground truth even on source distirbution images, it can result in noiser data and more difficult training. The choice should be validated experimentally for each model and dataset.